OpenStudy (anonymous):

What is the average force required of the brakes in a 1200 kg car in order for it to come to rest from 60kmh–1 in a distance of 150m?

OpenStudy (anonymous):

Given: m = mass = 1,200 kg vo = initial velocity = 60 kph = 16.667 m/sec vf = final velocity = 0 (it come to rest) s = distance = 150m a = acceleration/deceleration = ? F = force = ? vf^2 = vo^2 + 2as 0 = (16.667)^2 + (2*a*150) -300a = 277.778 a = 277.778/-300 a = -0.926 m/sec^2 (negative acceleration because it is coming to stop/rest) F = ma F = 1,200kg x 0.926 m/sec^2 F = 1,111.11 Newtons (answer)

OpenStudy (anonymous):

Thank you so much!

mathslover (mathslover):

first of all good work @nitinz570 I will request you to mention the link also: http://au.answers.yahoo.com/question/index?qid=20110509050436AAvB8tG

mathslover (mathslover):

From which you have taken the answer ..

OpenStudy (anonymous):

i have written myself....understand

mathslover (mathslover):

sorry nitin you are wrong.. that link contains exact ditto soln